Please show calculation how you solved the problem. The Web-based company Oh Baby! has a goal ofprocessing 95 percent of its orders on the same day they are received. If 485 out of the next 500 orders are processed on the same day, would this prove that they are exceeding their goal, using d = .025?(not sure of the symbol is a or d,copy it the from text book).
Answer by stanbon(57219)

Please show calculation how you solved the problem. The Web-based company Oh Baby! has a goal of processing 95 percent of its orders on the same day they are received. If 485 out of the next 500 orders are processed on the same day, would this prove that they are exceeding their goal, using alpha = .025
--------------------
Ho: p = 0.95
Ha: p > 0.95
-----------------
sample proportion: 485/500 = 0.97
---
Critical value for one-tail test with alpha = 2.5%: z = 1.96
Teststatic: z(0.97)=(0.97-0.95)/sqrt[0.95*0.05/500] = 2.052
p-value: P(z>2.052) = 0.02
--------------------------------
Conclusion: Since the teststatistic is in the "reject" interval,
reject Ho.
OR
Since the p-value is less than alpha, reject Ho.
--------------------------------
Cheers,
Stan H.